# Measuring deviations from spherical symmetry

### v1

@@ Appendix A Preliminary technical results
In this section we present and prove several preliminary results, which will be essential for the proofs of the statements in Section 3 and 4. Our first result provides some properties of the Fisher-von-Mises distribution.
@@
By equation (9) in Marco Di Marzio and Taylor, (2014), we have
@@
as $\kappa\to\infty$ and with the notation $a_{j}(p)=2^{j/2}\Gamma((p+j-1)/2)/\Gamma((p-1)/2)$ the result (5.7) follows. For a proof of (5.8) we note that the modified Bessel function in (2.11) satisfies $\mathcal{I}_{v}(\kappa)\sim e^{\kappa}/\sqrt{2\pi\kappa}$, as $\kappa\to\infty$ (see equation (9.6.18) on Page 376 of Abramowitz and Stegun, (1968)). Therefore, recalling the definition of the Fisher-von-Mises distribution in (2.10), we obtain
@@
and note that (see the proof of Theorem 3.1 in Marco Di Marzio and Taylor, (2014))
@@
(see the proof of Theorem 3.1 in Marco Di Marzio and Taylor,, 2014), which yields
@@ Appendix B Proof of main results
Then, if $hf_{0}$ is the density of a spherical symmetric distribution (for some density $h$ on $\mathbb{R}^{+}$), we obtain for the squared $L^{2}$-distance between $f$ and $hf_{0}$
@@ B.2 Martingale structure under spherical symmetry
By Lemma B.1, ${\cal M}_{n}^{2}$ is a sum of martingale difference and we can apply a central limit theorem for sums of martingale differences (Theorem 1 in Hall,, 1984) to prove the statement in Theorem 3.1. We have already shown that $H_{n}$ is symmetric, $\mathbb{E}\{H_{n}(Y_{1},Y_{2})|Y_{1}\}=0$ a.s., and by Proposition 3.1 it follows that $\mathbb{E}\{H_{n}^{2}(Y_{1},Y_{2})\}<\infty$ for each $n$. Therefore it is sufficient to verify that the condition
@@
$R_{21}$ $=\Big\|\int_{\mathbb{R}^{+}}\int_{\mathcal{\mathbb{}}{S}^{p-1}}h^{-1}c_{1}^{-1}(\kappa)K\Big(\frac{u_{2}-U_{1}}{h}\Big)L(\kappa v_{2}^{\top}V_{1})f(u_{2},v_{2})\mathrm{d}u_{2}\ \omega_{p-1}(\mathrm{d}v_{2})-f_{U,V}(U_{1},V_{1})\Big\|,$ (5.201)
@@ A Taylor expansion gives for the first term
$R_{21}$ $=\Big\|\int_{-1}^{1}\int_{\mathcal{\mathbb{}}{S}^{p-1}}c_{1}^{-1}(\kappa)K\left(u\right)L(\kappa v_{2}^{\top}V_{1})f(U_{1}+hu,v_{2})\mathrm{d}u\ \omega_{p-1}(\mathrm{d}v_{2})-f_{U,V}(U_{1},V_{1})\Big\|$ (5.204)
@@
$~~~~~~~~~~~~~~~~~~\times f(U_{1}+hu,V_{1}\cos\theta+\xi\sin\theta)(\sin\theta)^{p-2}\mathrm{d}u\ \mathrm{d}\theta\ \omega_{p-2}(\mathrm{d}\xi)-f_{U,V}(U_{1},V_{1})\Big\|$ (5.206)
@@
Therefore, by Donsker’s Theorem (see the discussion on page 225 - 226 in Van Der Vaart and Wellner,, 1996) it follows that

### v2

@@ Appendix A Preliminary technical results
In this section, we present and prove several preliminary results, which will be essential for the proofs of the statements in Section 3 and 4. Our first result provides some properties of the Fisher-von-Mises distribution.
@@
By equation (9) in DiMarzio03042014, we have
@@
as $\kappa\to\infty$ and with the notation $a_{j}(p)=2^{j/2}\Gamma((p+j-1)/2)/\Gamma((p-1)/2)$ the result (5.8) follows. For a proof of (5.9) we note that the modified Bessel function in (2.11) satisfies $\mathcal{I}_{v}(\kappa)\sim e^{\kappa}/\sqrt{2\pi\kappa}$, as $\kappa\to\infty$ (see equation (9.6.18) on Page 376 of abramowitz1968handbook). Therefore, recalling the definition of the Fisher-von-Mises distribution in (2.10), we obtain
@@
and note that (see the proof of Theorem 3.1 in DiMarzio03042014)
@@
(see the proof of Theorem 3.1 in DiMarzio03042014), which yields
@@ Appendix B Proof of main results
Then, if $hf_{0}$ is the density of a spherically symmetric distribution (for some density $h$ on $\mathbb{R}^{+}$), we obtain for the squared $L^{2}$-distance between $f$ and $hf_{0}$
@@ B.2 Martingale structure under spherical symmetry
By Lemma B.1, ${\cal M}_{n}^{2}$ is a sum of martingale difference and we can apply a central limit theorem for sums of martingale differences (Theorem 1 in HALL19841) to prove the statement in Theorem 3.1. We have already shown that $H_{n}$ is symmetric, $\mathbb{E}\{H_{n}(Y_{1},Y_{2})|Y_{1}\}=0$ a.s., and by Proposition 3.1 it follows that $\mathbb{E}\{H_{n}^{2}(Y_{1},Y_{2})\}<\infty$ for each $n$. Therefore it is sufficient to verify that the condition
@@
$R_{21}$ $=\Big\|\int_{\mathbb{R}^{+}}\int_{\mathcal{\mathbb{}}{S}^{p-1}}h^{-1}c_{1}^{-1}(\kappa)K\Big(\frac{u_{2}-U_{1}}{h}\Big)L(\kappa v_{2}^{\top}V_{1})f(u_{2},v_{2})\mathrm{d}u_{2}\ \omega_{p-1}(\mathrm{d}v_{2})-f(U_{1},V_{1})\Big\|,$ (5.202)
@@ A Taylor expansion gives for the first term
$R_{21}$ $=\Big\|\int_{-1}^{1}\int_{\mathcal{\mathbb{}}{S}^{p-1}}c_{1}^{-1}(\kappa)K\left(u\right)L(\kappa v_{2}^{\top}V_{1})f(U_{1}+hu,v_{2})\mathrm{d}u\ \omega_{p-1}(\mathrm{d}v_{2})-f(U_{1},V_{1})\Big\|$ (5.205)
@@
$~~~~~~~~~~~~~~~~~~\times f(U_{1}+hu,V_{1}\cos\theta+\xi\sin\theta)(\sin\theta)^{p-2}\mathrm{d}u\ \mathrm{d}\theta\ \omega_{p-2}(\mathrm{d}\xi)-f(U_{1},V_{1})\Big\|$ (5.207)
@@
Therefore, by Donsker’s Theorem (see the discussion on page 225 - 226 in van1996weak) it follows that